# HarmAug: Effective Data Augmentation for Knowledge Distillation of Safety Guard Models

### v2

@@ 3.2 Data Augmentation: HarmAug
To address this issue, we propose a simple prefix attack to bypass the safety guardrail of $p_{\texttt{LLM}}$. In addition to a set of $k$ exemplars $\{{\mathbf{x}}_{j_{1}},\ldots,{\mathbf{x}}_{j_{k}}\}$ randomly sampled from $\mathcal{D}$ and a prompt describing how to generate harmful instructions, such as “Make a single harmful instruction prompt that would elicit offensive content.”, we add an affirmative prefix of the LLM’s response to the prompt (e.g., “I have an idea for a prompt:”) as follows:
@@
Given the prompt with the affirmative prefix, denoted as ${\mathbf{z}}_{j}$, the LLM completes the response, i.e., $\hat{{\mathbf{x}}}_{j}\sim p_{\texttt{LLM}}(\cdot\mid{\mathbf{z}}_{j})$, leading to the sampling harmful instructions. We refer to our method as HarmAug. Empirically, we found that our prefix attack effectively bypasses the built-in guardrails of the LLM, allowing for the generation of harmful instructions (Sec. 4.4). This jailbreak vulnerability may be attributed to a weakness in the current RLHF process for safety alignment. Humans rarely respond with a refusal immediately following an affirmative answer to a request, and the LLM is supervised fine-tuned to replicate such human behavior before the RLHF process. As a result, the model is heavily biased towards generating refusal responses to harmful instructions but the model is rarely penalized for generating responses after an affirmative prefix during RLHF, despite the prompt being harmful.
After sampling synthetic harmful instructions, we utilize two different LLMs for generating responses to those synthetic harmful instructions. The first LLM generates a refusal, denoted as $\hat{{\mathbf{y}}}_{j1},$ to each harmful instruction $\hat{{\mathbf{x}}}_{j}$. Similarly, the second LLM, which is fine-tuned on few-shot adversarial examples, samples a harmful response $\hat{{\mathbf{y}}}_{j2}$ to each $\hat{{\mathbf{x}}}_{j}$. Then, the teacher $p_{\theta}$ labels each instruction-response pair:
@@
for $l\in\{1,2\}$, where $\mathbbm{1}$ is an indicator function and $\tau\in(0,1)$ is a threshold for the pair of sequences classified as harmful. Finally, we augment the training dataset with our synthetic dataset $\hat{\mathcal{D}}=\{(\hat{{\mathbf{x}}}_{j},\hat{{\mathbf{y}}}_{j1},c_{j1}),(\hat{{\mathbf{x}}}_{j},\hat{{\mathbf{y}}}_{j2},c_{j2})\}_{j=1}^{m}$ and train the small safety guard model $q_{\phi}$ with Eq. 1.
@@
For the training dataset $\mathcal{D}$, we use the train split of WildGuardMix (Han et al., 2024). We evaluate the safety guard models on four public benchmark datasets: OpenAI Moderation (OAI; Markov et al., 2023), ToxicChat (Lin et al., 2023), HarmBench (Mazeika et al., 2024), and the test split of WildGuardMix. The first two datasets are targeted for instruction classification (i.e., a response is always an empty sequence), while the others are designed for instruction-response pair classification.
@@
    GFN (Lee et al., 2024): This approach trains an LM with GFlowNet (GFN; Bengio et al., 2021) to sample harmful instructions proportional to the mixture of the harmful score distribution induced by the safety guard model $p_{\theta}$ and a reference language model’s likelihood. We augment the training $\mathcal{D}$ with instructions generated by the LM fine-tuned with GFlowNet and train DeBERTa on the augmented dataset. More details are provided in Sec. A.1.
@@
Following prior works (Inan et al., 2023; Han et al., 2024), we evaluate the safety guard models using F1 score and AUPRC. More details are provided in Appendix B.
@@ 4.1 Main Results
Table 1: We run experiments three times with different random seeds and report the average of F1 and AUPRC scores. The best results are bolded and the second-best are underlined. For results including standard deviations, please refer to  Table 8 in Appendix C.
@@
Llama-Guard-2 8B $\underline{0.7635}$ $0.8441$ $0.4233$ $0.4368$ $0.7777$ $0.8802$ $0.6585$ $0.7652$ $0.6557$ $0.7316$
Llama-Guard-3 8B 0.7884 $\underline{0.8750}$ $0.4859$ $0.4823$ $0.8445$ 0.8959 $0.6998$ $0.8127$ $0.7046$ $0.7665$
WildGuard [^1] We report “n/a” for AUPRC since the WildGuard library does not provide the probability of harmfulness. 7B $0.7268$ n/a $0.6547$ n/a 0.8596 n/a $0.7504$ n/a 0.7479 n/a
@@
DeBERTa + GFN 435M $0.6939$ $0.7793$ $0.6259$ $0.7191$ $\underline{0.8463}$ $\underline{0.8842}$ $0.7443$ 0.8376 0.72760.7276{0.7276}0.7276 $0.8050$
DeBERTa + HarmAug 435M $0.7236$ 0.8791 $0.6283$ 0.7553 $0.8331$ 0.88410.8841{0.8841}0.8841 0.7576 $0.8265$ $\underline{0.7357}$ 0.8362
@@
To study how our data augmentation method changes distribution of instructions, we cluster the prompts from the union of the original dataset $\mathcal{D}$ and our synthetic dataset $\hat{\mathcal{D}}$, and compare it against clustering with only the original dataset. We use Hugging Face’s text clustering library which embeds sentences with a language model and runs DBSCAN (Ester et al., 1996) for clustering. As shown in Fig. 3, our data augmentation significantly increases the number of clusters from 65 to 332. This suggests our data augmentation method, HarmAug, improves diversity of instructions in the training dataset. Generated instructions are presented in Table 11 of Appendix D.
@@
[…]
Background.
$R({\mathbf{x}})=\exp\left(\frac{1}{\beta}\mathbb{E}_{{\mathbf{y}}\sim p_{\texttt{target}}({\mathbf{y}}\mid{\mathbf{x}})}[\log p_{\theta}(c=1\mid{\mathbf{x}},{\mathbf{y}})]\right)+p_{\texttt{ref}}({\mathbf{x}})^{1/\gamma},$ (3)
where $\beta$ and $\gamma$ are positive constants that control the peakiness of the reward, $p_{\theta}$ is a safety guard model, and $p_{\texttt{ref}}$ is a reference language model to measure the likelihood of ${\mathbf{x}}$ to enforce the generation of natural sentences. Then the language model $p_{\psi}$ is trained with the following trajectory balance objective (Malkin et al., 2022):
Experimental setup.
$\frac{1}{5k}\sum_{i=1}^{k}\sum_{j=1}^{5}p_{\theta}(c=1\mid{\mathbf{x}}^{(i)},{\mathbf{y}}^{(j)}),\quad{\mathbf{x}}^{(i)}\overset{\mathrm{iid}}{\sim}p_{\psi}({\mathbf{x}}),\quad{\mathbf{y}}^{(j)}\overset{\mathrm{iid}}{\sim}p_{\psi}({\mathbf{y}}\mid{\mathbf{x}}^{(i)})$ (5)
which is referred to as “Test Reward” in Table 3. For diversity, following prior work (Hong et al., 2024), we calculate the average cosine distance between all possible pairs of the generated prompts. Please refer to Sec. A.2 for more details.
@@ Results.
Figure 4: (a): AUPRC score on various jailbreak attacks with our model (DeBERTa-large) and Llama-Guard-3. (b): Plot of AUPRC score on CipherChat as a function of wallclock time when fine-tuning the models.
Figure 5: After further fine-tuning DeBERTa and Llama-Guard-3 on CipherChat and WildGuardMix datasets, we report average F1 and AUPRC score of five runs on each dataset.
@@
We further fine-tune Llama-Guard-3 and DeBERTa-large trained with our HarmAug method on the CipherChat (Yuan et al., 2024) dataset, which comprises pairs of harmful instructions and responses encoded in ASCII for the purpose of jailbreak. To prevent catastrophic forgetting (McCloskey & Cohen, 1989), we sample a mini-batch from both the WildGuardMix and CipherChat datasets in every update step. We train the models using LoRA (Hu et al., 2022) for $200$ steps, with the rank set to $32$, a batch size of $8$, and a learning rate of $10^{-4}$. Finally, we evaluate the models by measuring F1 and AUPRC scores on both the WildGuardMix and CipherChat.
@@ 4.4 Ablations
In this section, we conduct a comprehensive ablation study of each component of our method to evaluate its effectiveness.
@@
To study the effectiveness of the prefix attack for generating harmful instructions, we remove the prefix “I have an idea for a prompt:” from the Prompt Format described in Sec. 3.2 and measure how often the LLM $p_{\texttt{LLM}}$ successfully generates instructions instead of refusing to do so. We sample $10,000$ instructions from $p_{\texttt{LLM}}$ and use a simple pattern matching classifier proposed by Zou et al. (2023) to evaluate whether the LLM refuses to generate instructions. As shown in Sec. 4.4, removing the prefix significantly degrades the success rate, which indicates the necessity of prefix attack for circumventing the safety alignment of the LLM.
Table 5: With different LLM backbones for sampling harmful instructions, we report the average F1 and AUPRC scores of three runs. For results including standard deviations, please refer to Table 9 in Appendix C.
@@
DeBERTa + Phi-3.5 3.8B $0.7230$ $0.8647$ $0.6073$ $0.7180$ $0.8337$ $0.8807$ $0.7543$ $0.8259$ $0.7295$ $0.8223$
DeBERTa + Mistral-0.3 7B $0.7317$ $0.8717$ $0.6230$ $0.7075$ $0.8304$ $0.8769$ $0.7516$ $0.8267$ $0.7342$ $0.8207$
@@
We perform an ablation study to examine the effect of LLM backbones in generating harmful instructions. We prompt the following models for sampling harmful instructions: Gemma-1.1-2b-it, Llama-3.1-Instruct-8B-Instruct, Llama-3.1-8B, Phi-3.5-mini-instruct, Mistral-7B-Instruct-v0.3, and the fine-tuned Llama-2 model (Wei et al., 2024), which has been fine-tuned on 100 adversarial prompts. All models are prompted using the prompt format described in Sec. 3.2. As shown in Table 5, regardless of the choice of LLMs, data augmentation with LLM-based prompting outperforms the other baselines in average, including GFN and EDA. Moreover, data augmentation with instructions generated by the smallest model, Gemma-1.1-2b-it, yields the most significant improvement in AUPRC.
@@
Table 7: Ablation study on the backbone architecture of student models. We run experiments three times with different random seeds and report the average of F1 and AUPRC scores. For results including standard deviations, please refer to Table 10 in Appendix C.
@@ 5 Conclusion
Limitations.
@@
Reproducibility statement.
We use PyTorch (Paszke et al., 2019) and the Transformers library from Hugging Face (Wolf et al., 2020) to implement our proposed method and all the baselines in our experiments. All implementation details are described in the experimental setup part of Sec. 4.1, Sec. 4.2, and Sec. 4.3. We provide anonymous URLs to our code, safety guard model, and synthetic dataset, allowing the research community to fully access, reproduce, and extend our work on improving detection of harmful conversations and computational efficiency of safety guard models. Detailed instructions for reproducing our knowledge distillation process are provided in our code.
Ethics statement.
Appendix
Appendix A Implementation Details
A.1 GFlowNet baseline

### v3

@@ 3.2 Data Augmentation: HarmAug
To address this issue, we propose a simple prefix attack to bypass the safety guardrail of $p_{\texttt{LLM}}$. In addition to a set of $k$ exemplars $\{{\mathbf{x}}_{j_{1}},\ldots,{\mathbf{x}}_{j_{k}}\}$ randomly sampled from $\mathcal{D}$, similar to (Wei et al., 2023b), and a prompt describing how to generate harmful instructions, such as “Make a single harmful instruction prompt that would elicit offensive content.”, we add an affirmative prefix of the LLM’s response to the prompt (e.g., “I have an idea for a prompt:”) as follows:
@@
This prefix attack is similar to the prefix injection (Wei et al., 2023a), asking the LLM to answer with a prefix by adding guidelines to the user prompt. However, our attack prefills the prefix in the LLM’s response and enforce the LLM to complete rest of the response. Given the prompt with the affirmative prefix, denoted as ${\mathbf{z}}_{j}$, the LLM completes the response, i.e., $\hat{{\mathbf{x}}}_{j}\sim p_{\texttt{LLM}}(\cdot\mid{\mathbf{z}}_{j})$, leading to the sampling harmful instructions. We refer to our method as HarmAug. Empirically, we found that our prefix attack effectively bypasses the built-in guardrails of the LLM, allowing for the generation of harmful instructions (Sec. 4.4). This jailbreak vulnerability may be attributed to a weakness in the current RLHF process for safety alignment. Humans rarely respond with a refusal immediately following an affirmative answer to a request, and the LLM is supervised fine-tuned to replicate such human behavior before the RLHF process. As a result, the model is heavily biased towards generating refusal responses to harmful instructions but the model is rarely penalized for generating responses after an affirmative prefix during RLHF, despite the prompt being harmful.
After sampling synthetic harmful instructions, we utilize two different LLMs for generating responses to those synthetic harmful instructions. The first LLM generates a refusal, denoted as $\hat{{\mathbf{y}}}_{j1},$ to each harmful instruction $\hat{{\mathbf{x}}}_{j}$. Similarly, the second LLM, which is fine-tuned on few-shot adversarial examples, samples a harmful response $\hat{{\mathbf{y}}}_{j2}$ to each $\hat{{\mathbf{x}}}_{j}$. Additionally, we pair the prompt with an empty sequence $\hat{\mathbf{y}}_{j3}$. The rationale for including the empty sequence is to train versatile safety guard models capable of handling both instruction classification and instruction-response pair classification tasks. Then, the teacher $p_{\theta}$ labels each instruction-response pair:
@@
for $l\in\{1,2,3\}$, where $\mathbbm{1}$ is an indicator function and $\tau\in(0,1)$ is a threshold for the pair of sequences classified as harmful. Finally, we augment the training dataset with our synthetic dataset $\hat{\mathcal{D}}=\{(\hat{{\mathbf{x}}}_{j},\hat{{\mathbf{y}}}_{jl},c_{jl})_{l=1}^{3}\}_{j=1}^{m}$ and train the small safety guard model $q_{\phi}$ with Eq. 1.
@@
For the training dataset $\mathcal{D}$, we use the train split of WildGuardMix (Han et al., 2024) combined with our synthetic dataset. We evaluate the safety guard models on four public benchmark datasets: OpenAI Moderation (OAI; Markov et al., 2023), ToxicChat (Lin et al., 2023), HarmBench (Mazeika et al., 2024), and the test split of WildGuardMix. The first two datasets are targeted for instruction classification (i.e., a response is always an empty sequence), while the others are designed for instruction-response pair classification.
@@
    GFN (Lee et al., 2024): This approach trains an LM with GFlowNet (GFN; Bengio et al., 2021) to sample harmful instructions proportional to the mixture of the harmful score distribution induced by the safety guard model $p_{\theta}$ and a reference language model’s likelihood. We augment the training $\mathcal{D}$ with instructions generated by the LM fine-tuned with GFlowNet and train DeBERTa on the augmented dataset. More details are provided in Sec. B.2.
@@
Following prior works (Inan et al., 2023; Han et al., 2024), we evaluate the safety guard models using F1 score and AUPRC. More details are provided in Appendix C.
@@ 4.1 Main Results
Table 1: We run experiments three times with different random seeds and report the average of F1 and AUPRC scores. The best results are bolded and the second-best are underlined. For results including standard deviations, please refer to Table 9 in Appendix D.
@@
Llama-Guard-2 8B 0.8139 $0.8824$ $0.4233$ $0.4368$ 0.8610 $0.8945$ $0.6870$ $0.7833$ $0.6963$ $0.7492$
Llama-Guard-3 8B ${\underline{0.8061}}$ 0.8869 $0.4859$ $0.4823$ 0.85510.8551{0.8551}0.8551 0.8999 0.68520.6852{0.6852}0.6852 0.81290.8129{0.8129}0.8129 0.70800.7080{0.7080}0.7080 0.77200.7720{0.7720}0.7720
WildGuard [^1] We report “n/a” for AUPRC since the WildGuard library does not provide the probability of harmfulness. 7B $0.7268$ n/a $0.6547$ n/a $\underline{0.8596}$ n/a $0.7504$ n/a 0.7479 n/a
@@
DeBERTa + GFN 435M $0.6939$ $0.7793$ $0.6259$ $0.7191$ 0.84630.8463{0.8463}0.8463 $\underline{0.8842}$ $0.7443$ 0.8376 0.72760.7276{0.7276}0.7276 $0.8050$
DeBERTa + HarmAug 435M $0.7236$ $\underline{0.8791}$ $0.6283$ 0.7553 $0.8331$ 0.88410.8841{0.8841}0.8841 0.7576 $0.8265$ $\underline{0.7357}$ 0.8362
@@
To study how our data augmentation method changes distribution of instructions, we cluster the prompts from the union of the original dataset $\mathcal{D}$ and our synthetic dataset $\hat{\mathcal{D}}$, and compare it against clustering with only the original dataset. We use Hugging Face’s text clustering library which embeds instructions with a language model and runs DBSCAN (Ester et al., 1996) for clustering. As shown in Fig. 3, our data augmentation significantly increases the number of clusters from 65 to 332. This suggests our data augmentation method, HarmAug, improves diversity of instructions in the training dataset. Generated instructions are presented in Table 12 of Appendix E.
[…]
Background.
$R({\mathbf{x}})=\exp\left(\frac{1}{\beta}\mathbb{E}_{{\mathbf{y}}\sim p_{\texttt{target}}({\mathbf{y}}\mid{\mathbf{x}})}[\log p_{\theta}(c=1\mid{\mathbf{x}},{\mathbf{y}})]\right)\cdot p_{\texttt{ref}}({\mathbf{x}})^{1/\gamma},$ (3)
where $\beta$ and $\gamma$ are positive constants that control the peakiness of the reward, $p_{\theta}$ is a safety guard model, and $p_{\texttt{ref}}$ is a reference language model to measure the likelihood of ${\mathbf{x}}$ to enforce the generation of natural sentences. Then the language model $p_{\psi}$ is trained to minimize the following trajectory balance objective (Malkin et al., 2022):
Experimental setup.
$\frac{1}{5k}\sum_{i=1}^{k}\sum_{j=1}^{5}p_{\theta}(c=1\mid{\mathbf{x}}^{(i)},{\mathbf{y}}^{(j)}),\quad{\mathbf{x}}^{(i)}\overset{\mathrm{iid}}{\sim}p_{\psi}({\mathbf{x}}),\quad{\mathbf{y}}^{(j)}\overset{\mathrm{iid}}{\sim}p_{\texttt{{target}}}({\mathbf{y}}\mid{\mathbf{x}}^{(i)})$ (5)
which is referred to as “Test Reward” in Table 3. For diversity, following prior work (Hong et al., 2024), we calculate the average cosine distance between all possible pairs of the generated prompts. Please refer to Sec. B.3 for more details.
@@ Results.
Figure 4: (a): Test AUPRC score on various jailbreak attacks with our model (DeBERTa-large) and Llama-Guard-3. (b): Plot of test AUPRC score on CipherChat as a function of wallclock time during fine-tuning.
Figure 5: After further fine-tuning DeBERTa and Llama-Guard-3 on CipherChat and WildGuardMix datasets, we report average test F1 and AUPRC score of five runs on each dataset.
@@
We further fine-tune Llama-Guard-3 and DeBERTa-large trained with our HarmAug method on the CipherChat (Yuan et al., 2024) dataset, which comprises 25 pairs of harmful instructions and responses encoded in ASCII for the purpose of jailbreak. To prevent catastrophic forgetting (McCloskey & Cohen, 1989), we sample a mini-batch from both the WildGuardMix and CipherChat datasets in every update step. We train the models using LoRA (Hu et al., 2022) for $200$ steps, with the rank set to $32$, a batch size of $8$, and a learning rate of $10^{-4}$. Finally, we evaluate the models by measuring F1 and AUPRC scores on both the test split of WildGuardMix and CipherChat.
@@ 4.4 Ablations
We conduct ablation studies of each component of our method to evaluate its effectiveness.
@@
To study the effectiveness of the prefix attack for generating harmful instructions, we remove the prefix “I have an idea for a prompt:” from the Prompt Format described in Sec. 3.2 and measure how often the LLM $p_{\texttt{LLM}}$ successfully generates instructions instead of refusing to do so. We sample $10,000$ instructions from $p_{\texttt{LLM}}$ and use a simple pattern matching classifier proposed by Zou et al. (2023) to evaluate whether the LLM refuses to generate instructions. As shown in Sec. 4.4, removing the prefix or replacing our prefix attack with the prefix injection attack proposed by Wei et al. (2023b), which instructs the LLM to begin its response with the affirmative prefix “Absolutely! Here’s ”, significantly degrades the success rate, which indicates the necessity of prefix attack for circumventing the safety alignment of the LLM.
Table 5: Different LLM backbones for sampling harmful instructions. We report the average F1 and AUPRC scores of three runs. Bold model names indicate LLMs used for our data augmentation method HarmAug. For results including standard deviations, please refer to Table 10 in Appendix D.
@@
DeBERTa + Phi-3.5 Instruct 3.8B $0.7230$ $0.8647$ $0.6073$ $0.7180$ $0.8337$ $0.8807$ $0.7543$ $0.8259$ $0.7295$ $0.8223$
DeBERTa + Mistral-0.3 Instruct 7B $0.7317$ $0.8717$ $0.6230$ $0.7075$ $0.8304$ $0.8769$ $0.7516$ $0.8267$ $0.7342$ $0.8207$
@@
We perform an ablation study to examine the effect of LLM backbones in generating harmful instructions. We prompt the following models for sampling harmful instructions: Gemma-1.1-2b-it, Llama-3.1-Instruct-8B-Instruct, Llama-3.1-8B, Phi-3.5-mini-instruct, Mistral-7B-Instruct-v0.3, and the fine-tuned Llama-2 model (Wei et al., 2024), which has been fine-tuned on 100 adversarial prompts. All models are prompted using the prompt format described in Sec. 3.2. As shown in Table 5, regardless of the choice of LLMs, data augmentation with LLM-based prompting outperforms the other baselines on average, including GFN and EDA. Moreover, data augmentation with instructions generated by the smallest model, Gemma-1.1-2b-it, yields the most significant improvement in AUPRC.
@@
Table 7: Ablation study on the backbone architecture of student models. We run experiments three times with different random seeds and report the average of F1 and AUPRC scores. For results including standard deviations, please refer to Table 11 in Appendix D.
@@ 5 Conclusion
Reproducibility statement
We use PyTorch (Paszke et al., 2019) and the Transformers library from Hugging Face (Wolf et al., 2020) to implement our proposed method and all the baselines in our experiments. All implementation details are described in the experimental setup part of Sec. 4.1, Sec. 4.2, and Sec. 4.3. We provide anonymous URLs to our code, safety guard model, and , allowing the research community to fully access, reproduce, and extend our work on improving detection of harmful conversations and computational efficiency of safety guard models. Detailed instructions for reproducing our knowledge distillation process are provided in our code.
Ethics Statement
Acknowledgment
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2020-II200153, Penetration Security Testing of ML Model Vulnerabilities and Defense), Institute for Information & communications Technology Promotion (IITP) grant funded by the Korea government (MSIT) (No.RS-2019-II190075 Artificial Intelligence Graduate School Program (KAIST)), Samsung Electronics (IO201214-08145-01), Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2022-II220713, Meta-learning Applicable to Real-world Problems), the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (NRF-2022R1A5A708390812), Institute of Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No.2022-0-00184, Development and Study of AI Technologies to Inexpensively Conform to Evolving Policy on Ethics), Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT(MSIT, Korea) & Gwangju Metropolitan City, Theori Inc., and JBin Project.
The authors would like to thank Heejun Lee for optimizing the inference speed of DeBERTa.
Appendix
Appendix A Limitations
@@
Appendix B Implementation Details
B.1 Model selection
We chose DeBERTa-v3-large based on the following three criteria. First, we prefer a bidirectional encoder over an autoregressive decoder-only model, as predicting the harmfulness of a prompt is a binary classification task rather than complex sequence generation. Second, among bidirectional encoders, we select the model based on its overall performance on general benchmark datasets, such as GLUE (Wang et al., 2024) and SQuAD (Rajpurkar et al., 2016). Finally, we choose the largest sub-billion-parameter model within the model family.
B.2 GFlowNet baseline